The air pressure inside the pressure cooker is 1.0 x 10⁵ Pa at the temperature of 28 °C.


What is the temperature, in kelvin needed to increase the pressure to 1.3 x 10⁵ Pa?


1
Expert's answer
2021-08-02T08:39:26-0400

The Charles's law says

"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

Hence,

"T_2=T_1\frac{P_2}{P_1}\\
=(28+273)\frac{1.3\times 10^5}{1.0\times 10^5}=391\:\rm K"

"t_2=T_2-273=391-273=118\:\rm ^{\circ}C"